Product Cost Method of Product Costing

Voice Com, Inc., uses the product cost method of applying the cost-plus approach to product pricing. The costs of producing and selling 5,080 units of cell phones are as follows:

Variable costs: Fixed costs:
Direct materials $66 per unit Factory overhead $199,000
Direct labor 40 Selling and admin. exp. 71,000
Factory overhead 26
Selling and admin. exp. 23
Total variable cost per unit $155 per unit

Voice Com desires a profit equal to a 16% rate of return on invested assets of $599,600.

a. Determine the amount of desired profit from the production and sale of 5,080 units of cell phones.
$

b. Determine the product cost per unit for the production of 5,080 of cell phones. If required, round your answer to nearest dollar.
$ per unit

c. Determine the product cost markup percentage (rounded to two decimal places) for cell phones.
%

d. Determine the selling price of cell phones. Round to the nearest dollar.

Total Cost $per unit
Markup per unit
Selling price $per unit
Answer #1

a) Desired profit form production = 5996000*16% = $95936

b) Product cost per unit = 66+40+26+(199000/5080) = 171 per unit

c) Markup percentage = (95936+71000/5080+23)/171 = 32.67%

d) Total cost = 171 per unit

Markup = (171*32.67%) = 56 per unit

Selling price = 171+56 = 227 per unit